OpenStudy (anonymous):

Solve the equation. 3x + 6 = 24 A. 2 B. 4 C. 6 D. 8

OpenStudy (anonymous):

i think its C.

OpenStudy (anonymous):

Let's go through it step by step ok?

OpenStudy (anonymous):

Do you know what the first thing you have to do is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok well we need to get x by itself, that is our ultimate goal. We are going to have to do something with the 6 right?

OpenStudy (anonymous):

umm change it to a -6

OpenStudy (kirbykirby):

3x + 6 = 24 3x = 24 - 6 3x = 18 x = 18/3 x = 6

OpenStudy (anonymous):

o well thanxs